# On the Approaching Geodesic Property via the Quotient Invariant

### v1

@@ 1. Introduction
If $D\Subset\mathbb{C}^{d}$ be a convex domain and $\lim_{z\to\partial D}q_{D}(z)=1$. Then $(D,K_{D})$ is a Gromov hyperbolic metric space that enjoys the visibility property. Moreover, the any two compactification $\overline{D}$, $\overline{D}^{G}$ and $\overline{D}^{H}$ are homeomorphic via the maps extending the identity.
@@ 2. Preliminaries and Technical Results
This proves the 3.5. ∎
Now, let us go back to the proof of the lemma. We have that $\gamma_{n}(K)\subset B(\phi_{n}(B_{R_{K}}),\varepsilon)$ for all $n\geq l_{R_{K},\varepsilon}$. Hence, in view of 3.5, we conclude that for all $n>N_{R,\varepsilon}:=\max\{l_{R_{K},\varepsilon},M_{R_{K},\varepsilon}\}$, for all $s,t\in K$ the following holds:
@@
Now invoking 3.5 and proceeding same way as Lemma 3.2 we conclude that $\psi_{n}\circ\tilde{\gamma}_{n}\to\tilde{\eta}$ uniformly on every compact subset of $\mathbb{R}$ and $\tilde{\eta}:\mathbb{R}\to\mathbb{B}^{d}$. This proves the lemma. ∎
@@
Let $K=[0,1]$ and $\varepsilon\in(0,\frac{c}{9})$. It follows from the proof of Lemma 3.3 and 3.7 that there exists $R_{K}>0$ and $\widetilde{N}_{R_{K},\varepsilon}=\max\{m_{R_{K},\varepsilon},l_{R_{K},\varepsilon}\}$ such that both $\gamma_{n}[0,1]$ and $\tilde{\gamma}_{n}[0,1]$ are inside $B(\phi_{n}(B_{R_{K}}),\varepsilon)$ for all $n\geq\widetilde{N}_{R_{K},\varepsilon}$. Therefore, in view of 3.5 we conclude that the following holds for all $t\in[0,1]$ and for all $n\geq\widetilde{N}_{R_{K},\varepsilon}$
@@ 4. Consequences of approaching geodesics
Here we present the proof of Theorem 1.8.

### v2

@@ 1. Introduction
If $D\Subset\mathbb{C}^{d}$ be a convex domain and $\lim_{z\to\partial D}q_{D}(z)=1$. Then $(D,K_{D})$ is a Gromov hyperbolic metric space that enjoys the visibility property. Moreover, any two compactification $\overline{D}$, $\overline{D}^{G}$ and $\overline{D}^{H}$ are homeomorphic via the maps extending the identity.
@@ 2. Preliminaries and Technical Results
This proves the Claim 3.5. ∎
Now, let us go back to the proof of the lemma. We have that $\gamma_{n}(K)\subset B(\phi_{n}(B_{R_{K}}),\varepsilon)$ for all $n\geq l_{R_{K},\varepsilon}$. Hence, in view of Claim 3.5, we conclude that for all $n>N_{R,\varepsilon}:=\max\{l_{R_{K},\varepsilon},M_{R_{K},\varepsilon}\}$, for all $s,t\in K$ the following holds:
@@
Now invoking Claim 3.5 and proceeding same way as Lemma 3.2 we conclude that $\psi_{n}\circ\tilde{\gamma}_{n}\to\tilde{\eta}$ uniformly on every compact subset of $\mathbb{R}$ and $\tilde{\eta}:\mathbb{R}\to\mathbb{B}^{d}$. This proves the lemma. ∎
@@
Let $K=[0,1]$ and $\varepsilon\in(0,\frac{c}{9})$. It follows from the proof of Lemma 3.3 and Claim 3.7 that there exists $R_{K}>0$ and $\widetilde{N}_{R_{K},\varepsilon}=\max\{m_{R_{K},\varepsilon},l_{R_{K},\varepsilon}\}$ such that both $\gamma_{n}[0,1]$ and $\tilde{\gamma}_{n}[0,1]$ are inside $B(\phi_{n}(B_{R_{K}}),\varepsilon)$ for all $n\geq\widetilde{N}_{R_{K},\varepsilon}$. Therefore, in view of Claim 3.5 we conclude that the following holds for all $t\in[0,1]$ and for all $n\geq\widetilde{N}_{R_{K},\varepsilon}$
@@ 4. Consequences of approaching geodesics
Here we present the proof of the Theorem   1.8.